# Humanoid Whole-Body Locomotion on Narrow Terrain via Dynamic Balance and Reinforcement Learning

### v1

@@ Abstract
Humans possess delicate dynamic balance mechanisms that enable them to maintain stability across diverse terrains and under extreme conditions. However, despite significant advances recently, existing locomotion algorithms for humanoid robots are still struggle to traverse extreme environments, especially in cases that lack external perception (e.g., vision or LiDAR). This is because current methods often rely on gait-based or perception-condition rewards, lacking effective mechanisms to handle unobservable obstacles and sudden balance loss. To address this challenge, we propose a novel whole-body locomotion algorithm based on dynamic balance and Reinforcement Learning (RL) that enables humanoid robots to traverse extreme terrains, particularly narrow pathways and unexpected obstacles, using only proprioception. Specifically, we introduce a dynamic balance mechanism by leveraging an extended measure of Zero-Moment Point (ZMP)-driven rewards and task-driven rewards in a whole-body actor-critic framework, aiming to achieve coordinated actions of the upper and lower limbs for robust locomotion. Experiments conducted on a full-sized Unitree H1-2 robot verify the ability of our method to maintain balance on extremely narrow terrains and under external disturbances, demonstrating its effectiveness in enhancing the robot’s adaptability to complex environments. The videos are given at https://whole-body-loco.github.io.
@@ I INTRODUCTION
In this paper, we propose a novel RL framework for whole-body locomotion in extreme scenarios, named Dynamic Balanced Humanoid Locomotion (DBHL). To enhance the locomotion policy’s ability to traverse complex terrains, particularly narrow pathways and sudden obstacles, we extend the concept of ZMP to non-planar surfaces, thereby forming a line of ZMPs. We then design a reward function for the RL policy that encourages the ZMP coordinates to be close to the center of the humanoid’s support polygon. This reward function is calculated using privileged information obtained from simulations, while the policy is learned solely based on proprioception via an asymmetric actor-critic framework. This design allows the policy to be deployed in real-world scenarios without relying on external perception. Within the RL framework, we train a whole-body control policy that leverages upper-body swings to assist dynamic balance. We introduce angular-momentum regularization and action-space noise to constrain undesired body rotation and action range. Finally, we integrate the ZMP-based reward with command-following and regularization rewards using a reward vectorization technique, where each reward term and value function are estimated independently, avoiding inaccurate estimation of small items in an accumulated reward.
In experiments, we evaluate DBHL under various challenging conditions, including walking on surfaces with unknown disturbances, stepping on stairs with different widths and heights, and traversing narrow slopes with varying widths and degrees. The results demonstrate that DBHL exhibits significantly better stability compared to other mainstream methods. We also analyze the role of the ZMP-reward through comparison, highlighting its crucial role in maintaining dynamic balance. Additionally, we conducted a detailed ablation study on other design elements. Our real-world experiments using the full-sized Unitree H1-2 robot illustrate the robot’s capability to move forward and backward in narrow terrain, handle disturbances such as pushes and trips, and showcase enhanced stability and adaptability in various extreme scenarios, as shown in Fig. 1.
@@
We construct a whole-body control framework with newly introduced techniques including reward vectorization, angular-momentum regularization and action-space noise.
@@ II-A Humanoid Locomotion
Research on humanoid locomotion can be traced back to the 1970s [14]. The fundamental idea for a locomotion controller is to decompose it into planning and tracking modules, where the planning module is responsible for generating desired trajectories, while the tracking module ensures that the robot follows these trajectories accurately [15, 16]. Methods such as Whole-Body Control (WBC) and Model Predictive Control (MPC) have achieved significant success in this domain [17, 18, 19]. However, these methods typically require precise modeling of the dynamics [20, 21], which poses substantial challenges for complex robot structures. In recent years, learning-based algorithms have emerged as a promising alternative to legged locomotion [1, 22] with efficient parallel simulation [8], significantly reducing the cost of interaction between the robot and the environment. Policies trained extensively in simulation environments can then be transferred to real robots [23]. In the field of quadrupedal locomotion, RL algorithms have demonstrated excellent performance in complex tasks such as complex-terrain walking [24, 25], gait control [26, 27], and even parkour [28, 29].
For RL-based humanoid locomotion, things become more difficult due to their limited support areas and high gravity. Meanwhile, classical control algorithms are also limited by the inaccurate modeling of the complex dynamical system. Recent approaches have proposed using RL algorithms for phase-based gait learning [2, 3], motor skill control [4, 5], and motion imitation [6, 7]. However, these methods still lags far behind that of humans in terms of dynamic balance under complex terrain (e.g., narrow paths) and extreme conditions (e.g., sudden disturbance). In our work, we address this problem by measuring the relationship between the ZMP and the humanoid support polygon in an RL framework. We note that a concurrent work uses foothold rewards to pass through narrow areas, while it relies on a LiDAR-based elevation map for real-world deployment [30]. In contrast, our method can traverse complex terrains only using proprioception without vision or LiDAR perception.
II-B Zero-moment Point
@@ III-A ZMP-based Dynamic Balance
To achieve dynamic balance, we design a ZMP-based reward and integrate it into the RL framework. With the standard definition, ZMP is the point on the ground where the horizontal component of the moment of the ground reaction forces becomes zero, assuming the ground is a horizontal plane. A necessary condition for the locomotion mechanism to be in dynamic equilibrium is that the candidate ZMP is inside the support polygon [31]. Traditional control methods achieve stable locomotion by planning the ZMP trajectory, and ensuring it stays within the support polygon. Furthermore, to handle non-planar surfaces, the concept of the Zero Moment Line (ZML) is introduced, which extends the ZMP concept to account for complex contact surfaces [35]. In DBHL, the ZMP-based reward is calculated by the distance between the ZML and the center of the support polygon formed by the robot’s feet, as shown in Fig. 3.
Then we give the calculation process as follows. Assuming that the ground reaction force is the only external force applied to the robot system, the moment of the ground reaction force about the origin of the world frame is given by
@@
where represents the linear momentum, the angular momentum, the total mass, the center of mass (CoM) position of the system, respectively, and . By the definition of ZMP, we have
@@
According to Eq. (4) and Eq. (5), ZMP at different heights lies at different locations, whose trajectory forms a line of ZMPs, called ZML [35]. The condition for stable locomotion is that the ZML must cross the support polygon.
Then, we need to estimate the support polygon of the humanoid robot to determine the relationship between ZML and the support. In general cases, the support polygon is determined by computing the convex hull of all contact points for scenarios with multiple contact points. However, the calculation becomes intricate when dealing with irregular contact surfaces, such as slopes or uneven terrain, where the contact points must be projected onto a horizontal plane, and the resulting convex hull is adjusted to account for the tilt [36]. Also, in movements, the support polygon should be updated to reflect the changing contact points. In DBHL, we can simplify the standard calculation because, in a RL framework, we only need to design a reward function to encourage the ZML to be close to the center of the support polygon, which can be achieved without strictly estimating the support polygon. Specifically, the geometric center of support polygon (denoted as ) is approximated by the center of supporting feet, as
@@
Based on the simplified , we design a computationally tractable reward function, denoted as , with the horizontal distance between and the ZML, as
where is a horizontal projection function that projects to the ZML. Intuitively, dynamic stability is guaranteed when is close to the ZML, and a smaller distance indicates better stability, as illustrated by Fig. 3.
@@ III-B1 Angular Momentum Regularization
It is introduced to minimize the undesired rotational motion during locomotion, thus improving coordination and improving resistance to external disturbances. As extensively discussed in prior work [37, 38], the motion of swinging legs generates significant angular momentum, which can disrupt whole-body motion. This effect can be mitigated through the proper use of upper limbs to counterbalance the momentum. We introduce a regularization reward based on the norm of the total angular momentum about the robot base position , which is defined as
where
where , , , , represent the position vector of the CoM, mass, inertia tensor, linear velocity, and angular velocity of the -th link, respectively.
@@ III-B2 Multiplicative Action Noise Injection
where . By applying large perturbations to actions with greater magnitudes, the policy is encouraged to favor small actions for the upper body joints, promoting cautious behavior when significant movements are required.
@@ III-C1 Task and Command
We developed three types of narrow terrain in policy training, including flat, stairs, and slopes. Meanwhile, a terrain-curriculum mechanism introduced in [8] is used in training with 20 difficulty levels, where the path width gradually decreases from 1.0 to 0.2 m, the slope gradient increases from 0 to 0.3, and the step height increases from 0 to 0.12 m. The robot is trained in a mixing of these narrow terrains, each containing 30% with an additional 10% planar terrain. For each episode, the linear velocity command is uniformly sampled in the range and , while the yaw velocity command is determined by where is the horizontal angle between positive X-axis direction and the orientation of the robot.
@@ III-C2 Asymmetric Actor-Critic Framework
III-C3 Reward and Value Vectorization
@@
The reward functions of DBHL are given in Table. I.
@@ III-C4 Symmetry Regularization
We introduce a symmetry loss term, inspired by [39], to enhance sample efficiency and promote more harmonious gaits. This loss leverages the symmetry of the robot’s motion with respect to the - plane, defined as
@@ III-C5 Domain Randomization
| Action Noise | , see Sec. III-B2 |
@@ IV EXPERIMENT
Q2: How does ZMP reward help with dynamic balance?
Q3: How does design choices (i.e., reward vectorization and action noise) influence training performance?
@@ IV-B2 Qualitative Comparison
In Fig. 7(a), we provide two example trajectories for a qualitative comparison. At s, an external push (m/s) is exerted on the robot. The trajectory without ZMP-based reward results in the robot losing balance, where the corresponding ZMP-distance, defined as in Eq. 7, becomes particularly large and eventually diverges, as shown in Fig. 7(b). In contrast, the policy incorporating ZMP-based reward successfully maintains the robot’s stability throughout the disturbance since the distance is constrained by maximizing the ZMP-reward in Eq. (7).
@@ IV-C Ablation Study
To address Q3, we perform ablation studies in Isaac Gym to investigate the effect of two key design choices: reward vectorization and multiplicative action noise, which aims to evaluate how each of these factors influences the performance of our method.
@@ IV-D2 Irregular Obstacles in Uneven Ground
Since the previous experiments focus solely on narrow terrains, we train an additional policy on irregular planar terrain. Then we evaluate the policy’s capabilities by incorporating a variety of irregular obstacles on flat terrain in the real world, as depicted in Figures 10 (d-f). This environment features acupressure plates, wooden blocks, planks, and stone-paved paths. In addition, the robot must contend with applied external forces while traversing these challenging conditions. The results shows our method can overcomes these complex terrain challenges.

### v2

@@ Abstract
Humans possess delicate dynamic balance mechanisms that enable them to maintain stability across diverse terrains and under extreme conditions. However, despite significant advances recently, existing locomotion algorithms for humanoid robots are still struggle to traverse extreme environments, especially in cases that lack external perception (e.g., vision or LiDAR). This is because current methods often rely on gait-based or perception-condition rewards, lacking effective mechanisms to handle unobservable obstacles and sudden balance loss. To address this challenge, we propose a novel whole-body locomotion algorithm based on dynamic balance and Reinforcement Learning (RL) that enables humanoid robots to traverse extreme terrains, particularly narrow pathways and unexpected obstacles, using only proprioception. Specifically, we introduce a dynamic balance mechanism by leveraging a novel Zero Moment Point (ZMP)-driven reward and task-driven rewards in a whole-body actor-critic framework, aiming to achieve coordinated actions of the upper and lower limbs for robust locomotion. Experiments conducted on a full-sized Unitree H1-2 robot verify the ability of our method to maintain balance on extremely narrow terrains and under external disturbances, demonstrating its effectiveness in enhancing the robot’s adaptability to complex environments. The videos are given at https://whole-body-loco.github.io.
@@ I INTRODUCTION
In this paper, we propose a novel RL framework for whole-body locomotion in extreme scenarios, named Dynamic Balanced Humanoid Locomotion (DBHL). To enhance the locomotion policy’s ability to traverse complex terrains, particularly narrow pathways and sudden obstacles, we extend the concept of ZMP to non-planar surfaces, thereby forming a line of ZMPs. We then design a reward function for the RL policy that encourages the ZMP coordinates to be close to the center of the humanoid’s support polygon. This reward function is calculated using privileged information obtained from simulations, while the policy is learned solely based on proprioception via an asymmetric actor-critic framework. This design allows the policy to be deployed in real-world scenarios without relying on external perception. Within the RL framework, we train a whole-body control policy that leverages upper-body swings to assist dynamic balance. We introduce angular momentum regularization and multiplicative action noise to constrain undesired body rotation and action range. Finally, we integrate the ZMP-based reward with command-following and regularization rewards using a reward vectorization technique, where each reward term is associated with a specific value function, avoiding inaccurate value estimation of small items in an accumulated reward.
In experiments, we evaluate DBHL under various challenging conditions, including walking on surfaces with unknown disturbances, stepping on stairs with different widths and heights, and traversing narrow slopes with varying widths and degrees. The results demonstrate that DBHL exhibits significantly better stability compared to other mainstream methods. We also analyze the role of the ZMP-reward through comparison, highlighting its crucial role in maintaining dynamic balance. Additionally, we conducted a detailed ablation study on other design elements. Our real-world experiments using the full-sized Unitree H1-2 robot illustrate the robot’s capability to traverse narrow terrains, handle disturbances such as pushes and trips, and showcase enhanced stability and adaptability in various extreme scenarios, as shown in Fig. 1.
@@
We construct a whole-body control framework with newly introduced techniques including reward vectorization, angular momentum regularization, and multiplicative action noise.
@@ II-A Humanoid Locomotion
Research on humanoid locomotion can be traced back to the 1970s [14]. The fundamental idea for a locomotion controller is to decompose it into planning and tracking modules, where the planning module is responsible for generating desired trajectories and the tracking module ensures that the robot follows these trajectories accurately [15, 16]. Methods such as Whole-Body Control (WBC) and Model Predictive Control (MPC) have achieved significant success in this domain [17, 18, 19]. However, these methods typically require precise modeling of the dynamics [20, 21], which poses substantial challenges for complex robot structures. In recent years, learning-based algorithms have emerged as a promising alternative to legged locomotion [1, 22] with efficient parallel simulation [8], significantly reducing the cost of interaction between the robot and the environment. Policies trained extensively in simulation environments can then be transferred to real robots [23]. In the field of quadrupedal locomotion, RL algorithms have demonstrated excellent performance in complex tasks such as complex-terrain walking [24, 25], gait control [26, 27], and even parkour [28, 29].
For RL-based humanoid locomotion, things become more difficult due to their limited support areas and high center of gravity. Meanwhile, classical control algorithms are also limited by the inaccurate modeling of the complex dynamical system. Recent approaches have proposed using RL algorithms for phase-based gait learning [2, 3], motor skill control [4, 5], and motion imitation [6, 7]. However, these methods still lags far behind that of humans in terms of dynamic balance under complex terrain (e.g., narrow paths) and extreme conditions (e.g., sudden disturbance). In our work, we address this problem by measuring the relationship between the ZMP and the humanoid support polygon in an RL framework. We note that a concurrent work uses foothold rewards to pass through narrow areas, while it relies on a LiDAR-based elevation map for real-world deployment [30]. In contrast, our method can traverse complex terrains only using proprioception without vision or LiDAR perception.
II-B Zero Moment Point
@@ III-A ZMP-based Dynamic Balance
To achieve dynamic balance, we design a ZMP-based reward and integrate it into the RL framework. The reward is calculated by the distance between a line of ZMPs, called Zero Moment Line (ZML) [35], and the center of the support polygon formed by the robot’s feet, as shown in Fig. 3.
Assuming that the ground reaction force is the only external force applied to the robot system, the moment of the ground reaction force about the origin of the world frame is given by
@@
where represents the linear momentum, the angular momentum about the origin, the total mass, the center of mass (CoM) position of the system, respectively, and . By the definition of ZMP, we have
@@
According to Eq. (4) and Eq. (5), ZMP at different heights lies at different locations, whose trajectory forms a line of ZMPs, called ZML [35].
The condition for stable locomotion is that the ZML must cross the support polygon. In planar surfaces, the support polygon is determined by computing the convex hull of all contact points for scenarios with multiple contact points. However, the calculation becomes intricate when dealing with irregular contact surfaces, such as slopes or uneven terrain, where contact points need to be projected onto a virtual plane [36]. In DBHL, we bypass explicit computation of the support polygon through reward design. In this way, the problem of dynamic balance is transformed into the design of a metric, which encourages the ZML to be close to the center of the support polygon.
Specifically, the geometric center of support polygon (denoted as ) is approximated by the center of supporting feet, as
@@
Based on the simplified , we design a computationally tractable reward function, denoted as , with the horizontal distance between and the ZML, called ZMP-distance, as
where is a projection function that projects to the ZML in the horizontal plane. Intuitively, dynamic stability is guaranteed when is close to the ZML, and a smaller distance indicates better stability, as shown in Fig. 3.
@@ III-B1 Angular Momentum Regularization
It is introduced to minimize the undesired rotational motion during locomotion, thus improving coordination and improving resistance to external disturbances. As extensively discussed in prior work [37, 38], the motion of swinging legs generates significant angular momentum, which can disrupt whole-body motion. This effect can be mitigated through the proper use of upper limbs to counterbalance the angular momentum. We introduce a regularization reward based on the norm of the total angular momentum about the robot base position, , which is defined as
where represents the mass of -th link, , , are the CoM position, linear velocity, and angular velocity of the -th link relative to the base position, respectively, and is the inertia tensor of the -th link about its CoM.
@@ III-B2 Multiplicative Action Noise Injection
where . By applying large perturbations to actions with greater magnitudes, the policy is encouraged to favor small actions for the upper body joints and to promote cautious behavior when significant movements are required.
@@ III-C1 Task and Command
We developed three types of narrow terrain in policy training, including flat, stairs, and slopes. The robot is trained in a mixing of these narrow terrains, each containing 30% with an additional 10% planar terrain. Meanwhile, a terrain-curriculum mechanism introduced in [8] is used in training with 20 difficulty levels, where the path width gradually decreases from 1.0 to 0.2m, the slope gradient increases from 0 to 0.3, and the step height increases from 0 to 0.12m. For each episode, the linear velocity command is uniformly sampled in the range and , while the yaw velocity command is determined by where is the horizontal angle between positive X-axis direction and the orientation of the robot.
@@ III-C2 Asymmetric Actor-Critic Framework
III-C3 Vectorization of Reward and Value
@@
The reward functions of DBHL are given in Table I.
@@ III-C4 Symmetry Regularization
Inspired by [39], we introduce a symmetry loss term to enhance sample efficiency and promote more harmonious gaits. This loss leverages the symmetry of the robot’s motion with respect to the - plane, defined as
@@ III-C5 Domain Randomization
| for each episode | |
| Action Noise | , see Sec. III-B2. |
@@ IV EXPERIMENT
Q2: How does the ZMP-based reward contribute to dynamic balance?
Q3: How do design choices (i.e., reward vectorization and action noise) influence training performance?
@@ IV-B2 Qualitative Comparison
In Fig. 7(a), we provide two example trajectories for a qualitative comparison. At s, an external push (m/s) is exerted on the robot. The trajectory without ZMP-based reward results in the robot losing balance, where the corresponding ZMP-distance, defined as in Eq. (7), becomes particularly large and eventually diverges, as shown in Fig. 7(b). In contrast, the policy incorporating ZMP-based reward successfully maintains the robot’s stability throughout the disturbance since the distance is constrained by maximizing the ZMP-reward in Eq. (7).
@@ IV-C Ablation Study
To address Q3, we perform ablation studies in Isaac Gym to investigate the effect of two key design choices: reward vectorization and multiplicative action noise.
@@ IV-D2 Irregular Obstacles in Uneven Ground
Since the previous experiments focus solely on narrow terrains, we train an additional policy on irregular planar terrain. Then we evaluate the policy’s capabilities by incorporating a variety of irregular obstacles on flat terrain in the real world, as depicted in Fig. 10 (d-f). This environment features acupressure plates, wooden blocks, planks, and stone-paved paths. In addition, the robot must contend with applied external forces while traversing challenging conditions. The results show DBHL can overcome these complex terrain challenges.